OpenStudy (anonymous):

Find an equation for the nth term of the sequence. -3, -12, -48, -192, ..

OpenStudy (freckles):

is there a common ratio or a common difference? this would be one of the first things I look for.

OpenStudy (anonymous):

common difference: 4

OpenStudy (freckles):

a common different is 4? but -12-(-3) isn't 4

OpenStudy (freckles):

do you mean common ratio -12/-3 is 4

OpenStudy (freckles):

since it has a common ratio it is a geometric sequence

OpenStudy (freckles):

the first term is -3 and the common ratio is 4 you have all you need to setup the formula for the nth term in the sequence

OpenStudy (freckles):

\[a_n=a_1 r ^{n-1} \text{ for integer } n \ge 2 \]

OpenStudy (anonymous):

a_n = -3 * 4^n-1

OpenStudy (freckles):

\[a_n=-3 \cdot 4^{n-1}\] is right